# PATCH - Psychometrics-AssisTed benCHmarking of Large Language Models: A Case Study of Mathematics Proficiency

### v1

@@ Abstract
Many existing benchmarks of large (multimodal) language models (LLMs) focus on measuring LLMs’ academic proficiency, often with also an interest in comparing model performance with human test takers. While these benchmarks have proven key to the development of LLMs, they suffer from several limitations, including questionable measurement quality (e.g., Do they measure what they are supposed to in a reliable way?), lack of quality assessment on the item level (e.g., Are some items more important or difficult than others?) and unclear human population reference (e.g., To whom can the model be compared?). In response to these challenges, we propose leveraging knowledge from psychometrics - a field dedicated to the measurement of latent variables like academic proficiency - into LLM benchmarking. We make three primary contributions. First, we introduce PATCH: a novel framework for Psychometrics-AssisTed benCHmarking of LLMs. PATCH addresses the aforementioned limitations, presenting a new direction for LLM benchmark research. Second, we implement PATCH by measuring GPT-4 and Gemini-Pro-Vision’s proficiency in 8th grade mathematics against 56 human populations. We show that adopting a psychometrics-based approach yields evaluation outcomes that diverge from those based on existing benchmarking practices. Third, we release 4 datasets to support measuring and comparing LLM proficiency in grade school mathematics and science against human populations.
@@ 1 Introduction
Despite the success of existing benchmarks in advancing LLM research, they are not without limitations. The first concern is measurement quality: Do these benchmarks measure what they are supposed to in a reliable way? Many benchmarks are created via crowd-sourced knowledge, by asking a convenience group of individuals (e.g., crowd workers, paper authors) to create new test items (e.g., GSM8K, DROP) or collecting them from (often undocumented) sources (e.g., websites, textbooks, school exams) (e.g., MATH, MMLU, ARC). Without domain expert input and rigorous testing of item quality, this approach often leads to undesirable outcomes, such as mismatch between a benchmark and its claimed measurement goal, missing information in a question, wrong answer keys, and low data annotation agreement (Nie et al., 2020). [^1] We avoid calling out specific datasets here, but a quick Internet search would reveal many blogs reporting large percentages of errors in existing LLM benchmarks.
@@
Second, current benchmarks do not account for differences across test items, such as item discriminativeness [^2] In psychometrics, the term “item discrimination” is used. However, given the ambiguity and negative connotation of “discrimination”, we adopt “discriminativeness” instead. and difficulty (see § 3.1). For example, consider three items A (easy), B (hard) and C (hard). While answering correctly to A and B would result in the same score as answering correctly to B and C, the latter (i.e., answering correctly to more difficult items) would imply higher proficiency. Furthermore, a benchmark that consists of only easy and hard items will fail to differentiate models with medium proficiency (i.e., low discriminativeness). Thus, without accounting for item differences, benchmarking results, especially model rankings, can be misleading.
@@
Third, while many benchmarks tried to compare LLMs against humans, the human population to be compared is unclear. For instance, human performance in MATH is based on the paper’s authors; in MMLU, crowd workers; in MATH, 6 university students in MATH. Using such convenience samples (with none to little information about sample characteristics), the resulting human performance is local to that specific sample and cannot be generalized to other human samples or specific populations.
@@
To address these challenges, we propose integrating insights from psychometrics — a field dedicated to the measurement of latent variables like cognitive abilities and academic proficiency — into the benchmarking process of LLMs. In particular, we draw on two research areas in psychometrics: item response theory (see § 3.1) and test development (see § 3.2 and 3.3). The former can help to estimate academic proficiency more accurately than common practice in LLM benchmarks (e.g., means, percentages, total scores). It can also provide diagnostic information about the quality of each test item. The latter, test development knowledge from psychometrics, can help to build high quality LLM benchmarks where comparison to specific human populations can be made.
@@
Our paper makes three primary contributions. First, we present PATCH: a novel framework for Psychometrics-AssisTed benCHmarking of LLMs, which addresses the aforementioned limitations of existing benchmarks. Second, we demonstrate the implementation of our framework by testing GPT-4 and Gemini’s proficiency in 8th grade mathematics using the released test items and data from Trends in International Mathematics and Science Study (TIMSS) 2011. We show empirically how adopting a psychometrics-based approach can lead to evaluation outcomes that diverge from those obtained through conventional benchmarking practices and are more informative, underscoring the potential of psychometrics to reshape the LLM benchmarking landscape. Third, we make our benchmark dataset and evaluation code based on TIMSS 2011 available to future researchers, along with three other math and science datasets based on TIMSS 2011 and 2008. We will encrypt them with a public key to avoid future data contamination, following the advice in Jacovi et al. (2023).
@@ 2 Related Work
We are not the first to think of leveraging psychometrics for LLM and more general NLP research. For instance, psychometric scales have been used to examine the psychological profiles of LLMs such as personality traits and motivations (Huang et al., 2024; Pellert et al., 2023; Dillion et al., 2023). The text in these scales can also be used to improve encoding and prediction of social science constructs like personality traits (Kreuter et al., 2022; Vu et al., 2020; Yang et al., 2021; Fang et al., 2023a). Psychometrics-based reliability and validity tests have also been proposed or/and used to assess the quality of model bias measures (Du et al., 2021; van der Wal et al., 2024), text embeddings (Fang et al., 2022), political stance detection (Sen et al., 2020), annotations (Amidei et al., 2020), user representations (Fang et al., 2023b), and general social science constructs (Birkenmaier et al., 2023).
The most closely related work to our paper is the use of IRT models in NLP for data splitting (Lalor et al., 2016), comparison of existing evaluation datasets and instances (e.g., difficulty, discriminativeness) (Sedoc & Ungar, 2020; Vania et al., 2021; Rodriguez et al., 2021; Lalor et al., 2018; Rodriguez et al., 2022), as well as identification of difficult instances from training dynamics (Lalor & Yu, 2020; Lalor et al., 2019). Our work distinguishes itself from these papers in two aspects. First, we do not apply IRT to existing LLM datasets/benchmarks. Instead, we introduce a framework for benchmarking LLMs by leveraging not only IRT but also test development knowledge from psychometrics. The goal of this framework is to generate new, high-quality benchmarks for LLMs that warrant valid comparison with human populations. Second, we demonstrate our framework with a mathematics proficiency test validated on human populations, and compare LLM performance with human performance. To the best our knowledge, we are the first to apply psychometrically validated (mathematics) proficiency tests to LLMs and make valid model/human comparisons.
@@ 3.1 Item Response Theory
IRT refers to a family of mathematical models that describe the functional relationship between responses to a test item, the test item’s characteristics (e.g., item difficulty and discriminativeness) and test taker’s standing on the latent construct being measured (e.g., proficiency) (AERA et al., 2014). Unlike classical test theory and current LLM benchmarks, which focus on the total or mean score of a test, IRT models takes into account the characteristics of both the items and the individuals being assessed, offering advantages like more accurate estimation of test takers’ proficiency, and item quality diagnostics. As such, IRT models have gained widespread adoption in various fields, including education, psychology, and healthcare, where precise measurement and assessment are crucial.
We describe below three fundamental IRT models suitable for different types of test items: the 3-parameter logistic (3PL) model for multiple choice items scored as either incorrect or correct, the 2-parameter logistic (2PL) model for open-ended response items scored as either incorrect or correct, as well as the generalized partial credit (GPC) model for open-ended response items scored as either incorrect, partially correct, or correct.
The 3PL model gives the probability that a test taker, whose proficiency is characterized by the latent variable $\theta$, will respond correctly to item $i$:
@@
where $x_{i}$ is the scored response to item $i$ ($1$ if correct and 00 if incorrect); $\theta$ is the proficiency of the test taker, where higher proficiency has a greater probability of responding correctly; $a_{i}$ is the slope parameter of item $i$, characterizing its discriminativeness (i.e., how well the item can tell test takers with higher $\theta$ from those with lower $\theta$) [^3] The number $1.7$ is a scaling parameter to preserve historical interpretation of parameter $a_{i}$ on the normal ogive scale (Camilli, 1994). Also applies to 2PL and GPC models.; $b_{i}$ is the location parameter of item $i$, characterizing its difficulty; $c_{i}$ is the lower asymptote parameter of item $i$, reflecting the chances of test takers with very low proficiency selecting the correct answer (i.e., guessing). Correspondingly, the probability of an incorrect response to item $i$ is: $P_{i,0}=P\left(x_{i}=0\mid\theta_{k},a_{i},b_{i},c_{i}\right)=1-P_{i,1}\left(\theta_{k}\right)$. The 2PL model has the same form as the 3PL model (Equation 1), except that the $c_{i}$ parameter is fixed at zero (i.e., no guessing).
@@
$P\left(x_{i}=l\mid\theta,a_{i},b_{i},d_{i,1},\cdots,d_{i,m_{i}-1}\right)=\frac{\exp\left(\sum_{v=0}^{l}1.7\cdot a_{i}\cdot\left(\theta-b_{i}+d_{i,v}\right)\right)}{\sum_{g=0}^{m_{i}-1}\exp\left(\sum_{v=0}^{g}1.7\cdot a_{i}\cdot\left(\theta-b_{i}+d_{i,v}\right)\right)}\\ \equiv P_{i,l}\left(\theta\right)$ (2)
@@
where $P_{i,l}\left(\theta\right)$ is of the form appropriate to the type of item (i.e., 3PL, 2PL or GPC), $m_{i}$ is equal to 2 for dichotomously scored items and 3 for polytomously scored items, and $u_{i,l}$ is an indicator variable defined as: $u_{i,l}=\left\{\begin{array}[]{l}1\text{ if response }x_{i}\text{ is in category }l\\ 0\text{ otherwise }\end{array}\right.$. This function can be viewed as a likelihood function to be maximized by the item parameters. With the estimated item parameters, $\theta$ can then be estimated (Reise & Revicki, 2014).
@@ 3.2 Test Development in Psychometrics
1. Construct and test need specification. 1. Test need (& construct) specification.
@@
Test development in psychometrics concerns the process of developing and implementing a test according to psychometric principles (Irwing & Hughes, 2018). Table 1 contrasts psychometric test development (based on Irwing & Hughes (2018)) with common LLM benchmarking procedures (based on Bowman & Dahl (2021); Raji et al. (2021)). What sets psychometric test development apart from typical LLM benchmark development is its focus on ensuring that the test matches a well-defined construct via expert-driven item generation, rigorous pilot testing, use of factor analysis and IRT models for item and test diagnostics, establishment of scoring and normalization standards, and testing on often representative samples of intended test takers. The result of this elaborate process is a high-quality test that can assess the construct of interest for the test takers in a valid and reliable way. Many large-scale assessments, such as PISA, TIMSS and PIRLS, confirm to such a process.
@@
We will use Proficiency in Grade School Mathematics (PGSM) as the construct of interest to further illustrate this process. In Step 1, the construct of interest and the test need are specified. For instance, what do we define PGSM? Is it based on a specific curriculum? What does existing literature say? Which education levels are we interested in? Is the test meant for comparison between students within a school, or between schools within a country? Such questions help us to clarify what we want to measure and how it can be measured.
@@
In Step 2, we make some necessary planning: How many test items? What kind of item format (e.g., multiple choice, short answer questions)? Will the test scores be standardized? How to assess the quality of test items? What are the desired psychometric properties of the test items (e.g., how discriminative and difficult should the items be?) and the test as a whole (e.g., internal consistency)? Will we pilot any test item? Will the test be computer- or paper-based? To sample test takers, what sampling frame and methodology should we use?
@@
In Step 3, we develop test items, which is an iterative procedure involving four steps: (a) construct refinement, where we further clarify the definition of PGSM (e.g., What content domains should be included: number, algebra, probability theory? Is proficiency only about knowing, or also about applying and reasoning?), (b) generate a pool of items with domain experts, (c) review the items for obvious misfit, errors and biases, and (d) pilot the items with an ideally representative sample of test takers. (e) With the responses from the pilot step, we can assess the psychometric properties of the test items with IRT and factor analysis (e.g., item discriminativeness; item difficulty; factor structure [^4] The correlational relationship between the test items intended to measure the construct of interest.). We iterate this procedure until we have a set of test items with acceptable psychometric properties. Then, in Step 4, we construct the PGSM test by specifying, for instance, which items to include (if not all), in which order, how many equivalent test versions, and what scoring instructions to use.
@@
In Step 5, the test gets implemented to the intended test takers, followed by Step 6: another round of quality analysis. If any item displays low quality characteristics (e.g., zero or negative discriminativeness), it will be left out of the final scoring. In Step 7, responses of the test takers are scored for each item, and the resulting item-level scores form the basis for estimating proficiency scores using IRT or simpler procedures like (weighted) sums. Normalizing the proficiency scores are also typical (e.g., a mean of 500 and a standard deviation of 100) to facilitate interpretations and comparisons. Finally, in Step 8, a technical manual is compiled, detailing all the results from Step 1-7, to facilitate correct re-use of the collected data, the test, as well as interpretation of test scores, among other purposes.

### v2

# PATCH! Psychometrics-AssisTed benCHmarking of Large Language Models: A Case Study of Proficiency in 8th Grade Mathematics

@@ Abstract
Many existing benchmarks of large (multimodal) language models (LLMs) focus on measuring LLMs’ academic proficiency, often with also an interest in comparing model performance with human test takers. While these benchmarks have proven key to the development of LLMs, they suffer from several limitations, including questionable measurement quality (e.g., Do they measure what they are supposed to in a reliable way?), lack of quality assessment on the item level (e.g., Are some items more important or difficult than others?) and unclear human population reference (e.g., To whom can the model be compared?). In response to these challenges, we propose leveraging knowledge from psychometrics - a field dedicated to the measurement of latent variables like academic proficiency - into LLM benchmarking. We make three primary contributions. First, we introduce PATCH: a novel framework for Psychometrics-AssisTed benCHmarking of LLMs. PATCH addresses the aforementioned limitations, presenting a new direction for LLM benchmark research. Second, we implement PATCH by measuring GPT-4 and Gemini-Pro-Vision’s proficiency in 8th grade mathematics against 56 human populations. We show that adopting a psychometrics-based approach yields evaluation outcomes that diverge from those based on existing benchmarking practices. Third, we release 4 high-quality datasets to support measuring and comparing LLM proficiency in grade school mathematics and science against human populations.
@@ 1 Introduction
Despite the success of existing benchmarks in advancing LLM research, they are not without limitations. The first concern is measurement quality: Do these benchmarks measure what they are supposed to in a reliable way? Many benchmarks are created via crowd-sourced knowledge, by asking a convenience group of individuals (e.g., crowd workers, paper authors) to create new test items (e.g., GSM8K, DROP) or collecting them from (often undocumented) sources (e.g., websites, textbooks, school exams) (e.g., MATH, MMLU, ARC). Without domain expert input and rigorous testing of item quality, undesirable outcomes can occur, including a mismatch between a benchmark and its claimed measurement goal, missing information in a question, wrong answer keys, and low data annotation agreement (Nie et al., 2020). [^1] We avoid calling out specific datasets here, but a quick Internet search would reveal many blogs reporting large percentages of errors in existing LLM benchmarks.
@@
Second, current benchmarks do not account for differences across test items, such as item discriminativeness [^2] In psychometrics, the term “item discrimination” is used. However, given the ambiguity and negative connotation of “discrimination”, we adopt “discriminativeness” instead. and difficulty (see Section 3.1). For example, consider three items A (easy), B (hard) and C (hard). While answering correctly to A and B would result in the same accuracy score as answering correctly to B and C, the latter (i.e., answering correctly to more difficult items) would imply higher proficiency. Furthermore, items that are too easy or too difficult (i.e., low discriminativeness) will fail to differentiate models of different proficiency levels. Thus, without accounting for item differences, benchmarking results, especially model rankings, can be misleading.
@@
Third, while many benchmarks are used to compare LLMs against humans, the human population to be compared is unclear (Tedeschi et al., 2023). For instance, human performance in MATH is based on the paper’s authors; in MMLU, crowd workers; in MATH, 6 university students. Using such convenience samples (with none to little information about sample characteristics), the resulting human performance is local to that specific sample and cannot be generalised to other human samples or specific populations.
@@
To address these challenges, we propose integrating insights from psychometrics - a field dedicated to the measurement of latent variables like cognitive abilities and academic proficiency - into LLM benchmarking processes. In particular, we draw on two research areas in psychometrics: item response theory (see Section 3.1) and test development (see Section 3.2 and 3.3). The former can help to estimate academic proficiency more accurately than common practice in LLM benchmarks (e.g., means, percentages, accuracy scores). It can also provide diagnostic information about the quality of each test item. The latter, test development knowledge, can help to build high quality LLM benchmarks where comparison to specific human populations can be made.
@@
Our paper makes three primary contributions. First, we present PATCH: a novel framework for Psychometrics-AssisTed benCHmarking of LLMs, which addresses the aforementioned limitations of existing benchmarks. Second, we demonstrate the implementation of PATCH by testing GPT-4 and Gemini’s proficiency in 8th grade mathematics using the released test items and data from Trends in International Mathematics and Science Study [^3] http://timssandpirls.bc.edu/timss2015/encyclopedia/ (TIMSS) 2011. We show empirically that a psychometrics-based approach can lead to evaluation outcomes that diverge from those obtained through conventional benchmarking practices and are more informative, underscoring the potential of psychometrics to reshape the LLM benchmarking landscape. Third, we make our benchmark dataset and evaluation code [^4] Available at https://github.com/fqixiang/patch_llm_benchmarking_with_psychometrics based on TIMSS 2011 available to future researchers, along with three other mathematics and science datasets based on TIMSS 2011 and 2008555Available at https://zenodo.org/records/12531906.
@@ 2 Related Work
We are not the first to propose leveraging psychometrics for research on LLMs and other areas in NLP. For instance, psychometric scales have been used to examine the psychological profiles of LLMs such as personality traits and motivations (Huang et al., 2024; Pellert et al., 2023; Dillion et al., 2023). The text in these scales can also be used to improve encoding and prediction of social science constructs like personality traits (Kreuter et al., 2022; Vu et al., 2020; Yang et al., 2021; Fang et al., 2023a). Psychometrics-based reliability and validity tests have also been proposed or/and used to assess the quality of NLP bias measures (Du et al., 2021; van der Wal et al., 2024), text embeddings (Fang et al., 2022), political stance detection (Sen et al., 2020), annotations (Amidei et al., 2020), user representations (Fang et al., 2023b), and general social science constructs (Birkenmaier et al., 2023).
The most closely related work to our paper is the use of IRT models in NLP for constructing more informative test datasets (Lalor et al., 2016), comparison of existing evaluation datasets and instances (e.g., difficulty, discriminativeness) (Sedoc & Ungar, 2020; Vania et al., 2021; Rodriguez et al., 2021; Lalor et al., 2018; Rodriguez et al., 2022), as well as identification of difficult instances from training dynamics (Lalor & Yu, 2020; Lalor et al., 2019). Our work distinguishes itself from these papers in two aspects. First, we do not apply IRT to existing LLM datasets/benchmarks. Instead, we introduce a framework for benchmarking LLMs by leveraging both IRT and test development knowledge from psychometrics. The goal of this framework is to generate new, high-quality benchmarks for LLMs that warrant valid comparison with human populations. Second, we demonstrate our framework with a mathematics proficiency test validated on 56 human populations, and compare LLM performance with human performance. To the best our knowledge, we are the first to apply psychometrically validated (mathematics) proficiency tests to LLMs and make valid model versus human comparison.
@@ 3.1 Item Response Theory
Item response theory (IRT) refers to a family of mathematical models that describe the functional relationship between responses to a test item, the test item’s characteristics (e.g., item difficulty and discriminativeness) and test taker’s standing on the latent construct being measured (e.g., proficiency) (AERA et al., 2014). Unlike classical test theory and current LLM benchmarks, which focus on the total or mean score of a test, IRT models takes into account the characteristics of both the items and the individuals being assessed, offering advantages like more accurate estimation of test takers’ proficiency, and item quality diagnostics. As such, IRT models have gained widespread adoption in various fields, including education, psychology, and healthcare, where precise measurement and assessment are crucial.
We describe below three fundamental IRT models suitable for different types of test items: the 3-parameter logistic (3PL) model for multiple choice items scored as either incorrect or correct, the 2-parameter logistic (2PL) model for open-ended response items scored as either incorrect or correct, as well as the generalised partial credit (GPC) model for open-ended response items scored as either incorrect, partially correct, or correct.
The 3PL model gives the probability that a test taker, whose proficiency is characterised by the latent variable $\theta$, will respond correctly to item $i$:
@@
where $x_{i}$ is the scored response to item $i$ ($1$ if correct and 00 if incorrect); $\theta$ is the proficiency of the test taker, where higher proficiency has a greater probability of responding correctly; $a_{i}$ is the slope parameter of item $i$, characterising its discriminativeness (i.e., how well the item can tell test takers with higher $\theta$ from those with lower $\theta$) [^6] The number $1.7$ is a scaling parameter to preserve historical interpretation of parameter $a_{i}$ on the normal ogive scale (Camilli, 1994). Also applies to 2PL and GPC models.; $b_{i}$ is the location parameter of item $i$, characterising its difficulty; $c_{i}$ is the lower asymptote parameter of item $i$, reflecting the chances of test takers with very low proficiency selecting the correct answer (i.e., guessing). Correspondingly, the probability of an incorrect response to item $i$ is: $P_{i,0}=P\left(x_{i}=0\mid\theta_{k},a_{i},b_{i},c_{i}\right)=1-P_{i,1}\left(\theta_{k}\right)$. The 2PL model has the same form as the 3PL model (Equation 1), except that the $c_{i}$ parameter is fixed at zero (i.e., no guessing).
@@
$P\left(x_{i}=l\mid\theta,a_{i},b_{i},d_{i,1},\cdots,d_{i,m_{i}-1}\right)$ $=\frac{\exp\left(\sum_{v=0}^{l}1.7\cdot a_{i}\cdot\left(\theta-b_{i}+d_{i,v}\right)\right)}{\sum_{g=0}^{m_{i}-1}\exp\left(\sum_{v=0}^{g}1.7\cdot a_{i}\cdot\left(\theta-b_{i}+d_{i,v}\right)\right)}$ (2)
$\equiv P_{i,l}\left(\theta\right)$
@@
where $P_{i,l}\left(\theta\right)$ is of the form appropriate to the type of item (i.e., 3PL, 2PL or GPC), $m_{i}$ is equal to 2 for dichotomously scored items and 3 for polytomously scored items, and $u_{i,l}$ is an indicator variable defined as:
$u_{i,l}=\left\{\begin{array}[]{l}1\text{ if response }x_{i}\text{ is in category }l\\ 0\text{ otherwise }\end{array}\right.$
This function can be viewed as a likelihood function to be maximised by the item parameters. With the estimated item parameters, $\theta$ can then be estimated (Reise & Revicki, 2014).
@@ 3.2 Test Development in Psychometrics
1. Construct and test need specification 1. (Construct and) test need specification
@@
Test development in psychometrics concerns the process of developing and implementing a test according to psychometric principles (Irwing & Hughes, 2018). Table 1 contrasts psychometric test development (based on Irwing & Hughes (2018)) with common LLM benchmarking procedures (based on Bowman et al. (2015); Raji et al. (2021)). What sets psychometric test development apart from typical LLM benchmark development is its focus on ensuring that the test matches a well-defined construct via expert-driven item generation, rigorous pilot testing, use of factor analysis and IRT models for item and test diagnostics, establishment of scoring and normalisation standards, and testing on representative samples of intended test takers. The result of this elaborate process is a high-quality test that can assess the construct of interest for the test takers in a valid and reliable way. Many large-scale assessments, such as PISA (Programme for International Student Assessment), TIMSS and PIRLS (Progress in International Reading Literacy Study), conform to such a process.
@@
We will use Proficiency in Grade School Mathematics (PGSM) as the construct of interest to further illustrate this process. In Step 1, the construct of interest and the test need are specified. For instance, how do we define PGSM? Is it based on a specific curriculum? What does existing literature say? Which education levels are we interested in? Is the test meant for comparison between students within a school, or between schools within a country? Such questions help us to clarify what we want to measure and how it can be measured.
@@
In Step 2, we make necessary planning: How many test items? What kind of item format (e.g., multiple choice, short answer questions)? Will the test scores be standardised? How to assess the quality of test items? What are the desired psychometric properties of the test items (e.g., how discriminative and difficult should the items be?) and the test as a whole (e.g., internal consistency)? Will we pilot any test item? Will the test be computer- or paper-based? To sample test takers, what kind of sampling frames and strategies should we use?
@@
In Step 3, we develop test items, which is an iterative procedure involving five steps: (a) construct refinement, where we further clarify the definition of PGSM (e.g., What content domains should be included: number, algebra, probability theory? Is proficiency only about knowing, or also about applying and reasoning?); (b) generate a pool of items with domain experts; (c) review the items for obvious misfit, errors and biases; (d) pilot the items with an ideally representative sample of test takers; (e) with the responses from the pilot step, we can assess the psychometric properties of the test items with IRT and factor analysis (e.g., item discriminativeness; item difficulty; factor structure [^7] Factor structure refers to the correlational relationships between the test items expected to capture the construct of interest.). We iterate this procedure until we have a set of test items with acceptable psychometric properties. Then, in Step 4, we construct the PGSM test by specifying, for instance, which items to include (if not all), in which order, how many equivalent test versions, and what scoring instructions to use.
@@
In Step 5, the test gets implemented to the intended test takers, followed by Step 6: another round of quality analysis. If any item displays low quality characteristics (e.g., zero or negative discriminativeness), it will be left out of the final scoring. In Step 7, responses of the test takers are scored for each item, and the resulting item-level scores form the basis for estimating proficiency scores using IRT or simpler procedures like (weighted) sums. Normalising the proficiency scores are also typical (e.g., a mean of 500 and a standard deviation of 100) to facilitate interpretations and comparisons. Finally, in Step 8, a technical manual is compiled, detailing all the results from Step 1–7, to facilitate correct re-use of the collected data, the test, as well as interpretation of test scores, among other purposes.